a truck is moving. when the truck is ahead of a car which is at rest, the truck having a speed of 50 m/s at that point accelerates at 1 m/s^2. three seconds later, the car starts and accelerates at 4m/s^2. when and where will the car overtake the truck? what are their velocities at that point?


1
Expert's answer
2021-05-09T13:04:02-0400

"s=v_0(t_0+t)+a_1(t_0+t)^2\/2=50(3+t)+(3+t)^2\/2"


"s=a_2t^2\/2=4t^2\/2=2t^2"


"50(3+t)+(3+t)^2/2=2t^2\to t=38.041\ (s)" . Answer


"s=2\cdot 38.041^2=2894\ (m)" . Answer


"v_1=50+1\cdot38.041=88.041\ (m/s)" . Answer


"v_2=4\cdot38.041=152.16 \ (m/s)" . Answer